The writing skills of high school seniors are different for boys and girls. that average writing test scores for girls are higher than those for boys. A model that best describes the writing test scores (w) for boys is w=-x+140. For girls a good model is w=1/4x+159. In both models x is the number of years after 1998. Assuming these models to be valid for the coming years, when will the aberage scores for girls be 49 points higher than for boys.
Sorry it's so long, but this one is really troubling me.
any help would be much appreciated!

Answer by stanbon(75887)

A model that best describes the writing test scores (w) for boys is w=-x+140. For girls a good model is w=1/4x+159. In both models x is the number of years after 1998. Assuming these models to be valid for the coming years, when will the average scores for girls be 49 points higher than for boys.
----
Solve:
girls ave - boys ave = 49
(1/4)x+159 - (-x+140) = 49
---
(1/4)x + 159 +x - 140 = 49
(5/4)x = -110
--
Comment: Something is wrong with the data you posted.
The answer is -88, which is probably wrong.
====
Cheers,
Stan H.